OpenStudy (anonymous):

a+b+c+d=360 then cosa+cosb+cosc+cosd=

hartnn (hartnn):

if a,b,c,d are angles in a cyclic quadrilateral, then a+c = 180 b+d = 180 and that expression will = 0 so is it given that a,b,c,d are angles in a cyclic quadrilateral?

OpenStudy (mathmale):

Bet that's an appropriate question, hartnn! But most of us on OpenStudy probably haven't heard the term "cyclic quadrilateral" before.

OpenStudy (anonymous):

http://openstudy.com/study#/updates/4fc8abc6e4b022f1e13024db I bet this is where hartnn got the idea from

OpenStudy (anonymous):

going back to the original question when a=b=c=d=90 cosa + cosb + cosc + cosd = 0 but when a=b=c=100 and d=60 cosa + cosb + cosc + cosd = -0.02... So I was unable to define a general solution for cosa+cosb+cosc+cosd when a+b+c+d=360

OpenStudy (mathmale):

Were I to experiment to determine whether there's a pattern here: I'd let each A, B, C and D equal 45 degrees and calculate cos A, cos B, cos C and cos D. Then I'd try other sets of angles whose sums are also 360 degrees. Not a fancy approach, but I'd expect it'd reveal some important facts/patterns here.